OpenStudy (anonymous):

Alaina has $72 and earns $6 each day. Joel has $48 and earns $8 each day. In how many days will Joel have as much money as Alaina?

OpenStudy (anonymous):

8x + 48 = 6x + 72

OpenStudy (anonymous):

So anyway set them equal to each other because the problem says they both have an equal amount of me on X number of days. Therefore you have to solve for X in this problem

OpenStudy (anonymous):

I see so when it says "How many days will Joel have as much money as Alaina that means how many days will they have equivalent (=) money right?

OpenStudy (anonymous):

What I would do is subtract 6x from both sides and subtract 48 from both sides you get 2x = 24 divide by two and x = 12

OpenStudy (anonymous):

x=12, right?

OpenStudy (anonymous):

Yes exactly

OpenStudy (anonymous):

Therefore in 12 days they will both have the same amount of money. you did it (;

OpenStudy (anonymous):

In my head

OpenStudy (anonymous):

So if i graphed it, what would be the point where they have same amount of money (12, ?)

OpenStudy (anonymous):

It would be helpful if you asked your second question :P

OpenStudy (anonymous):

according to my calculations if your asking how much money they would both have it would be (12, 144)

OpenStudy (anonymous):

Alaina has $72 and earns $6 each day, Joel has $48 and earns $8 each day. In How many days will joel have as much money as Alaina? A=72($)+6($/day)*D(day) J=48($)+8($/day)*D(day) A=J (Alaina and Joel must have the same ammount of money) (72+6D)($) = (48+8D)*($) (note that system units is $-money and simplifies at the end) 24=2D D=12

OpenStudy (anonymous):

We established that..:P